# Text2Data: Low-Resource Data Generation with Textual Control

### v1

@@ 5       Replace ${\mathbf{c}}$ with $\emptyset$ with probability $p_{uncond}$
4.3 Generalization bound of learning constraint
@@
Theorem 3 can be proved starting from the Bernstein’s inequality and the union bound inequality on squared zero-mean sub-Gaussian random variable. More detailed proof is in Appendix C. As a large amount of unlabeled data (i.e., $N$) is usually easy to obtain by either manual collection or simulation, $\epsilon_{N}$ is not large after taking logarithm on the number of model parameters (i.e., $|\Theta|$), even though it is usually much larger than $N$. Additionally, $\log|\Theta|$ is not significantly larger than $N_{p}$ so that $\epsilon_{N_{p}}$ should not be large as well. For instance, in our experiments, around 45,000 samples and 14 million model parameters for motion generation result in rather small $\epsilon_{N}$ and $\epsilon_{N_{p}}$. In practice, we use $\xi=\rho\cdot\inf_{\theta\in\Theta}{\mathcal{L}}_{1}(\theta)$ in Eq. 5 (i.e., $\hat{\xi}=\rho\cdot\inf_{\theta\in\hat{\Theta}}\hat{{\mathcal{L}}}_{1}(\theta)$ in Eq. 6) to relax the constraint, where $\rho$ is an hyperparameter to keep the constraint within the confidence interval.
5 Experiments
5.1 Datasets
We employ datasets from three modalities that may suffer from low-resource scenario.
Molecules. We extract 130,831 molecules from QM9 dataset (Ramakrishnan et al., 2014) with six molecular properties: polarizability ($\alpha$), highest occupied molecular orbital energy ($\epsilon_{\text{HOMO}}$), lowest unoccupied molecular orbital energy ($\epsilon_{\text{LUMO}}$), the energy difference between HOMO and LUMO ($\Delta_{\epsilon}$), dipole moment ($\mu$) and heat capacity at 298.15K ($C_{v}$).
Motions. We employ HumanML3D that contains textually re-annotating motions captured from AMASS (Mahmood et al., 2019) and HumanAct12 (Guo et al., 2020). It contains 14,616 motions annotated by 44,970 textual descriptions.
Time Series. We assemble 24 stocks from Yahoo Finance [^1] https://finance.yahoo.com/ (Appendix D) from their IPO date to July 8, 2023. We further tailor data to the length of 120 by slicing on the opening price for every 120 days, and scale by min-max normalization following Yoon et al. (2019). Totally, 210,964 time series are produced. We then extract their features including frequency, skewness, mean, variance, linearity (i.e., $R^{2}$), and the number of peaks via “tsfresh” in Python.
Each dataset is divided into training and testing sets at the ratio of $80\%$ and $20\%$, respectively. We curate each dataset to have varying proportions (i.e., $2\%,4\%,6\%,8\%,10\%,20\%,30\%,40\%$) of text labels. Details regarding constructing text descriptions for Molecules and Time Series are introduced in Appendix D.
5.2 Baseline models
@@
E(3) Equivariant Diffusion Model (EDM) (Hoogeboom et al., 2022).To handle Molecule dataset, we employ EDM as the baseline. EDM utilizes an equivariant network to denoise diffusion processes by concurrently processing both continuous (i.e., atom coordinates) and categorical data (i.e., atom types). The controllability on molecular properties is realized by the classifier-free diffusion guidance conditioned on the embedding of text descriptions, which is encoded by a pretrained T5 (Raffel et al., 2020) encoder (i.e., t5-large).
Motion Diffusion Model (MDM) (Tevet et al., 2023). We use MDM, a classifier-free diffusion model, for text-to-human motion generation. The text descriptions are embedded by a pretrained T5 encoder to guide the motion generation, providing the mechanism for controllability.
Generation diffusion for time series (DiffTS). To generate time series, we design the classifier-free diffusion model conditioned on text embeddings encoded from pretrained T5 encoder. We employ the backbone of (Ho & Salimans, 2022) by substituting image data to one-dimensional time series, and replacing the U-Net with one-dimensional convolutional neural network.
During implementation, we train the baselines on a specific proportion of labeled data. Text2Data is modified from the baseline model with a pretraining+finetuning strategy following Eq. 6. Additionally, we conduct ablation study where each baseline is still finetuned but without the constraint in Eq. 6 as a transfer learning-based baseline, or directly applied to augmented text-data pairs. More details are in Appendix E.
@@ 5.3 Evaluation metrics
Controllability. We compare the controllability between Text2Data and the baselines by the similarity between generated data and the ground truth. To assess the generated molecules, we follow (Hoogeboom et al., 2022) and train a classifier for each property to extract specific properties from generated data. Then, we calculate the Mean Absolute Error (MAE) between the extracted property and the ground truth. To assess the controllability of motion generation, we compute R precision and Multimodal distance that measure the relevancy of the generated motions to the input prompts (Guo et al., 2022b). To evaluate the controllability of time series generation, we extract properties via ”tsfresh” and compute the MAE between properties of generated data and that of ground truth. Additionally, we also visualize generated data according to the specified properties.
Generation quality. The evaluation of generation quality varies based on the modality. For molecular generation, we compute $-\log$ likelihood ($-\log p$(Hoogeboom et al., 2022) and validity of generated molecules (Jin et al., 2018), molecular stability and atom stability (Garcia Satorras et al., 2021) to evaluate their overall generation quality. For motion generation, we use FID score and Diversity following (Tevet et al., 2023). For time series generation, we follow (Yoon et al., 2019) and (Lim et al., 2023) by drawing t-SNE plots to visualize the overlap between generated data and the ground-truth. Better model tends to have a larger overlap, indicating more similar distribution.
@@
5.4 Comparisons on controllability
Figure 2 illustrates the MAE trend between the specific property of generated molecules and the intended one as the proportion of labeled training data rises. Text2Data achieves superior performance than EDM-finetune and EDM on all six properties by a remarkable margin. The results also indicate that certain properties, such as $\epsilon_{LUMO}$ and $Cv$, are more readily controllable. For these properties, the performance of the three models converges as the amount of labeled training data becomes sufficiently large (Figure 2). We further increase the proportion of available labels in the dataset up to $100\%$ and, as indicated in Appendix Table 3, the MAE keeps increasing when more labels are involved during training, and it gradually converges in the end. Appendix Table 4 also suggests that Text2Data surpasses the data augmentation-based method, which may suffer from the potentially poor alignment between text and data and the potential overfitting.
@@ 5.4 Comparisons on controllability
Additionally, we evaluate the controllability of Text2Data, along with its baseline comparisons, by utilizing MAE to measure the congruence between the property of generated data and the intended one within the Time Series dataset. As indicated in Table 2, Text2Data consistently excels over two baselines, DiffTS-finetune and DiffTS, across all three properties assessed during the experiments. Results of another three properties are presented in Appendix Table 5, which suggests the similar conclusion. Specifically, Text2Data and DiffTS-finetune both show a marked improvement over DiffTS in controlling frequency, variance, and skewness. They also exhibit a slight edge in controlling mean, number of peaks, and linearity. The enhanced performance of Text2Data correlates with its proficiency in alleviating the issue of catastrophic forgetting while maintaining a pursuit of controllability.
(a) Text2Data
(b) DiffTS-finetune
(c) DiffTS
5.5 Comparisons on generation quality
@@
We evaluate the overall quality of generating time series by making t-SNE plots of generated time series against ground truth. Substantial overlap between the generated time series and the ground truth suggests a closer distribution alignment, implying a better performance of Text2Data. As demonstrated in Figure 4, specifically, the red pattern represents the t-SNE of the ground-truth time series, whereas the blue pattern represents the t-SNE of the generated time series according to the same text description. Compared with DiffTS-finetune and DiffTS, Text2Data corresponds to the largest overlap between distributions of the generated and the ground-truth time series, suggesting its superior ability to precisely generate data according to the text description. The non-overlapping part may result from the diversity of generated time series or some other properties not labeled in the dataset so that not controlled by text description. The inferior performance of DiffTS stems from its training solely on labeled data, potentially leading to an incomplete understanding of the overall data distribution and a risk of overfitting due to limited size of labeled data. DiffTS may only partially capture the data distribution based on textual descriptions due to its susceptibility to catastrophic forgetting, which also heightens the risk of overfitting.
6 Conclusion
In this paper, we propose a novel constraint optimization-based framework, Text2Data, to improve the quality and controllability of text-to-data generation for various modalities in low-resource scenarios using diffusion models. Text2Data employs unlabeled data to capture the prevailing data distribution using an unsupervised diffusion model. It is then finetuned on text-labeled data, employing a novel constraint optimization-based learning objective to ensure controllability while reducing catastrophic forgetting. Comprehensive experiments are conducted on real-world datasets and our model shows consistently superior performance to recent baselines. While Text2Data is presented as a diffusion-based framework in this article, it can be seamlessly adapted to other generative models such as generative adversarial networks.
7 Ethical statement
We develop our method from publicly available QM9 (Ramakrishnan et al., 2014) and HumanML3D (Guo et al., 2022a) datasets, and stock data from public Yahoo Finance [^2] https://finance.yahoo.com/. It is important to note that, like other text-to-data models, our implementation will likely reflect the socioeconomic and entity biases inherent in datasets that we use. Additionally, although our method is designed for controllable data generation from text, we are not able to control the prompt that user inputs, which may contain improper contents.
@@ A.2 Reverse process
A.3 Derivation of learning objective
@@ Appendix B Lexicographic optimization
The model is firstly trained on unlabeled data, and then finetuned on labeled data under the framework of lexicographic optimization (Gong & Liu, 2021). Lexicographic optimization is used to balance the optimization of the learning objective (i.e., ^$\hat{{\mathcal{L}}}_{2}(\theta)$) and its constraint (i.e., ^$\hat{{\mathcal{L}}}_{1}^{\prime}(\theta)$). Specifically, lexicographic optimization leverage a dynamic gradient descent (Gong & Liu, 2021):
@@ Appendix C Proof of Theorem 1
The square of a zero-mean sub-Gaussian random variable with parameter $\sigma^{2}$ is a sub-exponential random variable with parameter $(4\sqrt{2}\sigma^{2},4\sigma^{2})$ (Honorio & Jaakkola, 2014).
@@ Appendix D More details regarding dataset
D.1 Stocks employed to form time series data
@@
D.2 Constructing text descriptions
@@ Appendix E Implementation details
E.1 Details of training process
We leverage a cross-attention layer to enhance the alignment between data and text embeddings (Ma et al., 2023). Specifically, we maximize the cosine-similarity between text embeddings obtained from the pretrained LLMs encoder and data embeddings diffused from original data, where the last layer of LLMs encoder is finetuned during training. Baselines models are trained for 3,000 epoches. Their finetuned version and Text2Data are both pretrained for 1,000 epoches and finetuned for another 2,000 epoches. All experiments are conduced on A100 hardware.
@@ E.1 Details of training process
E.2 Score functions
For time series generation, we use the same score function as in (Rasul et al., 2021) that consists of conditional 1-dim dilated ConvNets with residual connections adapted from the WaveNet (Oord et al., 2016) and Diffwave (Kong et al., 2020). For molecule generation, we employ the same score function from EDM (Hoogeboom et al., 2022) which is a Multilayer Perceptron. For motion generation, we also leverage the same score function in (Tevet et al., 2023) as a straightforward transformer.
@@ Appendix F Additional results
Proportion (%) $\alpha$ $\epsilon_{HOMO}$ Text2Data EDM-finetune EDM-finetune-unlabel EDM-DA EDM Text2Data EDM-finetune EDM-finetune-unlabel EDM-DA EDM 2 0.66 0.84 0.67 0.80 0.83 0.58 0.92 0.58 0.95 1.00 4 0.62 0.74 0.65 0.69 0.69 0.56 0.84 0.58 0.93 1.07 6 0.44 0.68 0.56 0.67 0.69 0.44 0.92 0.50 0.90 1.00 8 0.52 0.67 0.54 0.66 0.69 0.47 0.90 0.50 0.88 0.95 10 0.34 0.66 0.46 0.61 0.75 0.33 0.89 0.43 0.88 0.92
@@
Table 6 shows the results of evaluating the generation quality of Text2Data and baseline comparisons on Molecule dataset based on $-\log p$ and atom validity. Table 7 shows the results of evaluating the generation quality of Text2Data and baseline comparisons on Molecule dataset based on molecular and atom stability.

### v2

@@ 5       Replace ${\mathbf{c}}$ with $\emptyset$ with probability $p_{uncond}$
Generalization bound of learning constraint
@@
Theorem 3 can be proved starting from the Bernstein’s inequality and the union bound inequality on squared zero-mean sub-Gaussian random variable. More detailed proof is in Appendix. As a large amount of unlabeled data (i.e., $N$) is usually easy to obtain by either manual collection or simulation, $\epsilon_{N}$ is not large after taking logarithm on the number of model parameters (i.e., $|\Theta|$), even though it is usually much larger than $N$. Additionally, $\log|\Theta|$ is not significantly larger than $N_{p}$ so that $\epsilon_{N_{p}}$ should not be large as well. For instance, in our experiments, around 45,000 samples and 14 million model parameters for motion generation result in rather small $\epsilon_{N}$ and $\epsilon_{N_{p}}$. In practice, we use $\xi=\rho\cdot\inf_{\theta\in\Theta}{\mathcal{L}}_{1}(\theta)$ in Eq. 5 (i.e., $\hat{\xi}=\rho\cdot\inf_{\theta\in\hat{\Theta}}\hat{{\mathcal{L}}}_{1}(\theta)$ in Eq. 6) to relax the constraint, where $\rho$ is an hyperparameter to keep the constraint within the confidence interval.
Experiments
Datasets
We employ datasets from three modalities that may suffer from low-resource scenario. Details are in Appendix.
Molecules. We extract 130,831 molecules from QM9 dataset (Ramakrishnan et al. 2014) with six molecular properties: polarizability ($\alpha$), highest occupied molecular orbital energy ($\epsilon_{\text{HOMO}}$), lowest unoccupied molecular orbital energy ($\epsilon_{\text{LUMO}}$), the energy difference between HOMO and LUMO ($\Delta_{\epsilon}$), dipole moment ($\mu$) and heat capacity at 298.15K ($C_{v}$).
Motions. We employ HumanML3D that contains textually re-annotating motions captured from AMASS (Mahmood et al. 2019) and HumanAct12 (Guo et al. 2020). It contains 14,616 motions annotated by 44,970 textual descriptions.
Time Series. We assemble 24 stocks from Yahoo Finance from their IPO date to July 8, 2023. We tailor data to the length of 120 by slicing on the opening price for every 120 days, and scale by min-max normalization following Yoon, Jarrett, and Van der Schaar (2019). Totally, 210,964 time series are produced. We extract their features including frequency, skewness, mean, variance, linearity (i.e., $R^{2}$), and the number of peaks via “tsfresh” in Python.
Baseline models
@@
E(3) Equivariant Diffusion Model (EDM) (Hoogeboom et al. 2022).To handle Molecule dataset, we employ EDM as the baseline. EDM utilizes an equivariant network to denoise diffusion processes by concurrently processing both continuous (i.e., atom coordinates) and categorical data (i.e., atom types). The controllability on molecular properties is realized by the classifier-free diffusion guidance conditioned on the embedding of text descriptions, which is encoded by a pretrained T5 (Raffel et al. 2020) encoder (i.e., t5-large).
Motion Diffusion Model (MDM) (Tevet et al. 2023). We use MDM, a classifier-free diffusion model, for text-to-human motion generation. The text descriptions are embedded by a pretrained T5 encoder to guide the motion generation, providing the mechanism for controllability.
Generation diffusion for time series (DiffTS). To generate time series, we design the classifier-free diffusion model conditioned on text embeddings encoded from pretrained T5 encoder. We employ the backbone of (Ho and Salimans 2022) by substituting image data to one-dimensional time series, and replacing the U-Net with one-dimensional convolutional neural network.
For implementation, we train baselines on specific proportions of labeled data. Text2Data is modified from a pretraining+finetuning strategy following Eq. 6. Additionally, we conduct ablation study where each baseline is still finetuned but without the constraint in Eq. 6 as a transfer learning-based baseline, or directly applied to augmented text-data pairs. Besides, we adapt from (You et al. 2024) to form a semi-supervised framework, by training a classifier to predict molecular properties and generate pseudo labels for unlabelled molecules. More details are in Appendix.
@@
Controllability. We compare the controllability between Text2Data and the baselines by the similarity between generated data and the ground truth. To assess the generated molecules, we follow (Hoogeboom et al. 2022) and train a classifier for each property to extract specific properties from generated data. Then, we calculate the Mean Absolute Error (MAE) between the extracted property and the ground truth. To assess the controllability of motion generation, we compute R precision and Multimodal distance that measure the relevancy of the generated motions to the input prompts (Guo et al. 2022a). To evaluate the controllability of time series generation, we extract properties via ”tsfresh” and compute the MAE between properties of generated data and that of ground truth. Additionally, we also visualize generated data according to the specified properties.
Generation quality. Generation quality varies based on the modality. For molecular generation, we compute $-\log$ likelihood ($-\log p$(Hoogeboom et al. 2022) and validity of generated molecules (Jin, Barzilay, and Jaakkola 2018), molecular stability and atom stability (Garcia Satorras et al. 2021) to evaluate their overall generation quality. For motion generation, we use FID score and Diversity following (Tevet et al. 2023). For time series generation, we follow (Yoon, Jarrett, and Van der Schaar 2019) and (Lim et al. 2023) by drawing t-SNE plots to visualize the overlap between generated data and the ground-truth. Better model tends to have a larger overlap, indicating more similar distribution.
@@
Comparisons on controllability
Figure 2 illustrates the MAE trend between properties of generated molecules and the intended one as the proportion of labeled training data rises. Text2Data achieves superior performance than EDM-finetune and EDM on all properties by a remarkable margin. The results also indicate that certain properties, such as $\epsilon_{LUMO}$ and $Cv$, are more readily controllable. For these properties, the performance of the three models converges as the amount of labeled training data becomes sufficiently large (Figure 2). We further increase the proportion of available labels in the dataset up to $100\%$ and, as indicated in Appendix Table 3, the MAE keeps increasing when more labels are involved during training, and it gradually converges in the end. Appendix Table 4 also suggests that Text2Data surpasses the data augmentation-based method, which may suffer from the potentially poor alignment between text and data and the potential overfitting.
@@
Additionally, we evaluate controllability of Text2Data, along with its baseline comparisons, utilizing MAE to measure the congruence between the property of generated data and the intended one within the Time Series dataset. As indicated in Table 2, Text2Data consistently excels over two baselines, DiffTS-finetune and DiffTS, across all three properties assessed. Results of another three properties are presented in Appendix Table 5, which suggests the similar conclusion. Specifically, Text2Data and DiffTS-finetune both show a marked improvement over DiffTS in controlling frequency, variance, and skewness. They also exhibit a slight edge in controlling mean, number of peaks, and linearity. The enhanced performance of Text2Data correlates with its proficiency in alleviating the issue of catastrophic forgetting while maintaining a pursuit of controllability.
Comparisons on generation quality
@@
We evaluate the overall quality of generating time series by making t-SNE plots of generated time series against ground truth. Substantial overlap between the generated time series and the ground truth suggests a closer distribution alignment, implying a better performance of Text2Data. As demonstrated in Appendix Figure 4, specifically, the red pattern represents the t-SNE of the ground-truth time series, whereas the blue pattern represents the t-SNE of the generated time series according to the same text description. Compared with DiffTS-finetune and DiffTS, Text2Data corresponds to the largest overlap between distributions of the generated and the ground-truth time series, suggesting its superior ability to precisely generate data according to the text description. The non-overlapping part may result from the diversity of generated time series or some other properties not labeled in the dataset so that not controlled by text description. The inferior performance of DiffTS stems from its training solely on labeled data, potentially leading to an incomplete understanding of the overall data distribution and a risk of overfitting due to limited size of labeled data. DiffTS may only partially capture the data distribution based on textual descriptions due to its susceptibility to catastrophic forgetting, which also heightens the risk of overfitting.
Conclusion
We propose Text2Data to improve quality and property control of text-to-data generation for various modalities in low-resource scenarios using diffusion models. Text2Data uses unlabeled data to capture the prevailing data distribution via unsupervised diffusion model. It is then finetuned on text-labeled data, with a novel constraint optimization-based learning objective to ensure controllability while reducing catastrophic forgetting. Experiments show consistently superior performance to recent baselines. While Text2Data is presented as a diffusion-based framework in this article, it can be seamlessly adapted to other generative models.
@@ Appendix A Derivation of diffusion model
Derivation of learning objective
@@ Appendix B Lexicographic optimization
The model is firstly trained on unlabeled data, and then finetuned on labeled data under the framework of lexicographic optimization (Gong and Liu 2021). Lexicographic optimization is used to balance the optimization of the learning objective (i.e., ^$\hat{{\mathcal{L}}}_{2}(\theta)$) and its constraint (i.e., ^$\hat{{\mathcal{L}}}_{1}^{\prime}(\theta)$). Specifically, lexicographic optimization leverage a dynamic gradient descent (Gong and Liu 2021):
@@ Appendix C Proof of Theorem 1
The square of a zero-mean sub-Gaussian random variable with parameter $\sigma^{2}$ is a sub-exponential random variable with parameter $(4\sqrt{2}\sigma^{2},4\sigma^{2})$ (Honorio and Jaakkola 2014).
@@ Appendix D More details regarding dataset
Stocks employed to form time series data
@@
Splitting training and testing sets
Each dataset is divided into training and testing sets at the ratio of $80\%$ and $20\%$, respectively. We curate each dataset to have varying proportions (i.e., $2\%,4\%,6\%,8\%,10\%,20\%,30\%,40\%$) of text labels.
Constructing text descriptions
@@ Appendix E Implementation details
Details of training process
We leverage a cross-attention layer to enhance the alignment between data and text embeddings (Ma et al. 2023). Specifically, we maximize the cosine-similarity between text embeddings obtained from the pretrained LLMs encoder and data embeddings diffused from original data, where the last layer of LLMs encoder is finetuned during training. Baselines models are trained for 3,000 epoches. Their finetuned version and Text2Data are both pretrained for 1,000 epoches and finetuned for another 2,000 epoches. All experiments are conduced on A100 hardware.
@@
Score functions
For time series generation, we use the same score function as in (Rasul et al. 2021) that consists of conditional 1-dim dilated ConvNets with residual connections adapted from the WaveNet (Oord et al. 2016) and Diffwave (Kong et al. 2020). For molecule generation, we employ the same score function from EDM (Hoogeboom et al. 2022) which is a Multilayer Perceptron. For motion generation, we also leverage the same score function in (Tevet et al. 2023) as a straightforward transformer.
@@ Appendix F Additional results
Proportion (%) $\alpha$ $\epsilon_{HOMO}$ Text2Data EDM-finetune EDM-finetune-unlabel EDM-DA EDM-semi EDM Text2Data EDM-finetune EDM-finetune-unlabel EDM-DA EDM 2 0.66 0.84 0.67 0.80 0.87 0.83 0.58 0.92 0.58 0.95 1.00 4 0.62 0.74 0.65 0.69 0.86 0.69 0.56 0.84 0.58 0.93 1.07 6 0.44 0.68 0.56 0.67 0.78 0.69 0.44 0.92 0.50 0.90 1.00 8 0.52 0.67 0.54 0.66 0.74 0.69 0.47 0.90 0.50 0.88 0.95 10 0.34 0.66 0.46 0.61 0.74 0.75 0.33 0.89 0.43 0.88 0.92
@@
Table 6 shows the results of evaluating the generation quality of Text2Data and baseline comparisons on Molecule dataset based on $-\log p$ and atom validity.
Table 7 shows the results of evaluating the generation quality of Text2Data and baseline comparisons on Molecule dataset based on molecular and atom stability.
Figure 4 shows the t-SNE plot on time series data generated by Text2Data, DiffTS-finetune model and DiffTS, where Text2Data generated time series are more aligned with the ground truth indicated by their overlap.
(a) Text2Data
(b) DiffTS-finetune
(c) DiffTS
Appendix G Ethical statement
We develop our method from publicly available QM9 (Ramakrishnan et al. 2014) and HumanML3D (Guo et al. 2022b) datasets, and stock data from public Yahoo Finance. It is important to note that, like other text-to-data models, our implementation will likely reflect the socioeconomic and entity biases inherent in datasets that we use. Additionally, although our method is designed for controllable data generation from text, we are not able to control the prompt that user inputs, which may contain improper contents.